North Eastern Locomotive Preservation Group

  • Increase font size
  • Default font size
  • Decrease font size
Home
Q6 Blog

A record of events, activities and our work with the Q6



Saturday 20 February 2016

Print

Andy reports another good day, with progress made thanks to Bryan and the three JVs, Terry, Peter and Ian.

The pistons had arrived back so the piston ring grooves were cleaned out and measured so some new rings can be made.

The junior volunteers have removed some more wet sand from the left sand box after putting some of their training into practice, making a scoop to reach the sand.

Some of the lubricator pipes went back onto the loco and the right front axle box new pipes were fitted. We will only be able to do the first two axles at the moment as others will require the wheel sets to be removed, to get access to the top of the boxes.

The cylinder cover copper rings have been annealed.

The side rod bush clearances have now all been checked and recorded and the split pins renewed.

Terry has checked all the lubrication trimmings and oil pots. This also included removing any water and contaminated oil.

The safety valves will be taken to Hopetown on Monday evening for a start to be made on overhauling them, with Neal’s assistance, .

The arch formers have been put out ready for collection to take to Carnforth for the K1 to get its new arch.

Ian Pearson is going to look after the Wednesday work party with the following jobs:

  • Fit left leading axle box new pipes, use right as an example.
  • Continue with all cab valve maintenance.
  • Continue cleaning loco.
  • Continue with sand box and re-connect any linkages/pipes as required.
  • Clean round steam reverser and nip up all glands, fill locking cylinder with oil.
  • Horn stay bolt holes (marked ) to ream out and machine new bolt to fit hole.
Last Updated on Sunday, 20 March 2016 22:35
 

Wednesday 17 February 2016

Print

Another wet and cold day, as it poured down from start to finish, but thankfully not the expected snow. Because of the bad weather forecast for the day, the JVs postponed their attendance to Saturday, but Derek, Jon, Mike Oliver and myself, led by Ian, spent another happy day with the Q6. Fortunately the pump was behaving itself as the water was coming in as fast as it was getting it out.

The piston valves were completed, with the covers left off, but with three badly worn studs still needing attention on the left hand valve housing. The horn stay bolts were checked and the two leading bolts on the left hand leading horn stay were marked up as needing attention. All the others appeared sound and rang true when struck with the hammer.

Derek finished cleaning and oiling the rods, and Mike continued with his work on the cab gauges - thanks for the packing for them (and for the valves) which enabled progress to be made.

A start was made on the sanders. The left hand cab sandbox was emptied of as much sand as our arms would reach with a plastic tub as a scoop: if someone could bring a long handled ladle on Saturday that would help clear the bottom five inches or so. The deeper into the sandbox, the wetter the sand, but we resisted using it to make a model of the Q6!

Instead, it is in a large plastic bag next to the workshop wall. The rear left hand sand pipe was removed in the hope that we would be able to empty the sand box from below, but the flap which moves to allow the sand to drop through was firmly stuck in the closed position.

Disconnecting the rodding mechanism allowed the operation of both sandboxes on the middle of the running plate to be checked, and they were working fine. So attention needs to focus on the left hand cab sandbox and freeing up the handle, rodding and flap, which may mean dismantling the whole mechanism from below. The hole in the frame through which part of the mechanism protrudes is also fouling the movement of the external lever and may need angle grinding to trim some metal off to allow the lever to move freely.

Last Updated on Sunday, 20 March 2016 22:28
 

Saturday 13 February 2016

Print

Only Ian Pearson and James Pearcy at Deviation Shed on Saturday, but a good day, albeit very cold and wet. The pump in the pit was working well with the mesh filtering out the remaining dirt from the pit and enabling the water to get away.

The slide bars arrived about 9.40 and were off loaded by Barnie, the duty shed fitter, and put into Deviation shed using the tele handler. Many thanks to him for his help. The L/H top and bottom slide bars were then fitted to the locomotive using the hydraulic lifting table. This great piece of kit meant the job was completed by lunchtime. After lunch, James measured up the 8 spots on the slide bars required by Neal and will pass this information on to him. A start was then made on the piston valves. The R/H piston valve was released from its grip on the valve rod, which had caused problems on Wednesday, and the L/H valve was then successfully fitted using strops and the pull lift. The L/H valve was repacked with the follower still to fit. The R/H valve is still to repack, for which some more 16mm packing might be required. Andy is going to speak to Neal about this on Monday night, as well as the separate packing requested by Mike Oliver for the cab gauges from last Wednesday's working party.

So some significant progress.

Jobs for Wednesday:

  • Finish off the sliders for the mesh in the pit to protect the pump from rubbish.
  • Check sanders work. The front ones might, but the back ones definitely do not. Will have to empty the sand boxes and investigate/repair.
  • Finish off piston valves but leave off the front covers.
  • Side rod bush clearances to check with feeler gauges and record sizes. Will have to move loco to get right leading and intermediate (check the cable for the lights is not trapped under loco!).
  • Remove the pipes from leading axle, from oil box to axle box, mark up & put into storage with the new flexible pipes(should be three and one with Neal to get some fittings made). There will be three ferrules in the top of the axle boxes to remove (right inner one already removed as sample to get fittings made). These must not be lost in case we have to revert back to the original set up. Look in top of axle boxes and remove any water from side cavities.
  • Check all horn stay bolts: left leading and intermediate reported as needing attention by the MPD. Mark up with yellow paint any that are loose. New bolts are on the bench behind the grinder, holes will require reaming.
Last Updated on Sunday, 20 March 2016 22:28
 

Wednesday 10 February 2016

Print

A day of steady if unspectacular progress. Joined by Terry Newman, who prepared the gaskets for the safety valves and whistle, the working party managed to insert the right hand valve, but the packing has yet to be fitted to the valve spindle. During the cleaning of the valve ports and the studs, three studs on the left hand side valve housing were marked up with badly worn threads which will need attention. Mike Oliver continued his sterling work on the cab gauges, and Derek cleaned the left hand side rods. The refurbished baffle plate was also delivered and is now in the workshop, whilst the slide bars had been collected during the week and taken away to Darlington for machining.

 

Saturday 6 February 2016

Print

Andy reports that there was a really good turn out on Saturday, with eight volunteers and four JVs, so lots of work was achieved. The pistons, with the rods attached went to Sheffield last Tuesday to get the rods ground true. As no-one from NELPG was present, these were loaded by the MPD staff, so many thanks are due to all those who were involved. Hopefully the pistons will be back in the next two weeks, and then M-Machine can make some new piston rings and packings for the glands.

Peter and Trevor were once again cleaning and getting items ready for some paint. Once the cleaning of the steam heat pipe was completed, Peter painted it under the loco and tender before moving on to the steam chest casting and cylinder covers. Trevor helped get the whistle off and cut material for the JVs to use during their training session with Chris Parrish.

Neal and James did a survey of the left cross head and slide bars, in the light of comments on previous experience from Richard Pearson. After some deliberation, it was decided to remove them and ask M-Machine if they could machine the slide bars and put some white metal onto the edges of the slippers. Hopefully, we will have them back for next Saturday to start rebuilding that cylinder assembly.

Ian has done a marvellous job getting the boiler ready for the insurance exam: he has also checked that enough door joint gaskets are in stock to box the boiler up when the time comes.

Bryan and Steve were looking after our JVs. They had a really interesting day with instruction from Chris Parrish and were also having a go at painting, making a mesh to stop rubbish getting into the pump's sump, checking the side rod bush clearances using feeler gauges, and fitting the slacker pipe and valve in the cab.

Hopefully we have now come to the point with the maintenance where we are starting to turn the corner and putting things back together. While the locomotive certainly looks a lot better now we have some newly painted items, there is still a way to go yet.

Jobs for Wednesday:-

  • Continue with cab valves including whistle.
  • Make new gaskets for safety valves and whistle.
  • Clean valve ports of carbon and wire brush studs.
  • Fit piston valves and covers.
  • Fit packing to valve spindles.
Last Updated on Sunday, 20 March 2016 22:35
 


Page 43 of 46